OpenStudy (anonymous):

correctly factor 9x^2-60xy+100y^2

OpenStudy (anonymous):

can i wrongly factorize it for u? ;p

OpenStudy (anonymous):

-__- that wouldn't really be helpful, thanks.

OpenStudy (anonymous):

haha ok ill be helpful then :P

OpenStudy (anonymous):

I've given up. I don't know what to do..

OpenStudy (anonymous):

ill talk it through with u dont worry :) u know how to facotorize right?

OpenStudy (anonymous):

factorize*

OpenStudy (anonymous):

yes I know how to factorize. the -60xy is throwing me off though.

OpenStudy (anonymous):

and thanks! :)

OpenStudy (anonymous):

i always start off like this (x + )(x + )

OpenStudy (anonymous):

since the first two must multiply to get 9x^2, wat would u put to be the coefficients of the two x's?

OpenStudy (anonymous):

2? since it's two x's right or no?

OpenStudy (anonymous):

well 2x*2x = 4x^2 u want them to = 9x^2 so try again :)

OpenStudy (anonymous):

3x*3x=9x^2

OpenStudy (anonymous):

i agree, let try that and see if we get it and if we dont then we can try x*9x = 9x^2 So we got, (3x + )(3x + ) and according to 9x^2-60xy+100y^2, the last two numbers must multiply to get 100y^2. but how do u think we will get the y's?

OpenStudy (anonymous):

10y*10y would equal out to 100y^2 so I suppose 10y*10y

OpenStudy (anonymous):

ok then we'll try that =) (3x + 10y)(3x + 10y) ok now expand it and see if u get 9x^2-60xy+100y^2. if not u can try changing it from +10y to negative 10y and hope that works :P

OpenStudy (anonymous):

Alright thank you soo much for all of your help! :)

OpenStudy (anonymous):

wait! did it expand to get 9x^2-60xy+100y^2?????

OpenStudy (anonymous):

I got it figured out, I made it work.

OpenStudy (anonymous):

just making sure u got (3x - 10y)(3x - 10y) cause this is the answer not(3x + 10y)(3x + 10y) :)

OpenStudy (anonymous):

or u can say (3x - 10y)^2 :)